# Eyes Closed, Safety On: Protecting Multimodal LLMs via Image-to-Text Transformation

### v1

@@ Abstract
Multimodal large language models (MLLMs) have shown impressive reasoning abilities, which, however, are also more vulnerable to jailbreak attacks than their LLM predecessors. Although still capable of detecting unsafe responses, we observe that safety mechanisms of the pre-aligned LLMs in MLLMs can be easily bypassed due to the introduction of image features. To construct robust MLLMs, we propose ECSO (Eyes Closed, Safety On), a novel training-free protecting approach that exploits the inherent safety awareness of MLLMs, and generates safer responses via adaptively transforming unsafe images into texts to activate intrinsic safety mechanism of pre-aligned LLMs in MLLMs. Experiments on five state-of-the-art (SoTA) MLLMs demonstrate that our ECSO enhances model safety significantly (e.g., a 37.6% improvement on the MM-SafetyBench (SD+OCR), and 71.3% on VLSafe for the LLaVA-1.5-7B), while consistently maintaining utility results on common MLLM benchmarks. Furthermore, we show that ECSO can be used as a data engine to generate supervised-finetuning (SFT) data for MLLM alignment without extra human intervention.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) [14, 3, 62, 9, 22, 64] have attracted significant attention for their remarkable multimodal capabilities. Building upon Large Language Models (LLMs) [55, 27, 53, 12], they are aligned with a pre-trained visual encoder using text-image datasets, empowering LLMs to conduct conversations with image inputs. Despite the accomplishments, MLLMs encounter the challenges when inheriting the safety mechanism of their LLM predecessors. As later discussed in Sec. 3, although built upon LLMs well-aligned with the human morals and values [12, 56], MLLMs can be easily induced to generate the unethical content with the introduction of image inputs [68, 35, 44].
To protect MLLMs, one can repeat the training-based alignment strategies of LLMs on MLLMs, including Supervised Finetuning (SFT) [59, 32, 49, 7] and Reinforcement Learning from Human Feedback (RLHF) [43, 13, 46]. However, they require meticulous design of red-teaming queries that can successfully induce LLMs to generate harmful responses, and become more challenging when image inputs involved [44, 68]. Therefore, our question arises, “How can we, to the greatest extent, transfer the pre-aligned safety mechanisms of LLMs to MLLMs?”
In this paper, we start by conducting a throughout analysis about the safety assessment ability of MLLMs, where we observe that despite their susceptibility to the malicious queries, (i) MLLMs exhibit clear awareness of unsafe content in their own responses [7]. (ii) The safety mechanism of pre-aligned LLMs persists in MLLMs, but is “suppressed” by the introduction of image features. However, this can be restored simply by removing the images. Building upon these insights, we propose ECSO (Eyes Closed, Safety On), a novel training-free MLLM protection strategy exploiting the intrinsic safety mechanisms of pre-aligned LLMs. When presented with an image input with a user query, ECSO first leverages the safety awareness of MLLMs to assess safety of their own responses in a post-hoc manner. Once unsafe initial responses detected, ECSO converts the image inputs into texts via the query-aware image-to-text (I2T) transformation, and degrades MLLMs into (text-only) LLMs. Safe response generation without images is then performed to restore the safety mechanism of pre-aligned LLMs. Experiments on five MLLM benchmarks demonstrate that our ECSO can significantly enhance the model safety without sacrificing the utility performance on common MLLM benchmarks. Moreover, we show that ECSO can be adopted as a data engine to generate SFT data to align MLLMs without extra human intervention.
@@
    We demonstrate that MLLMs, although susceptible to jailbreaking attacks, can detect unsafe content in their own responses and also inherit the safety mechanisms from pre-aligned LLMs, although inadvertently suppressed.
@@
    ECSO significantly enhances the safety of five SoTA MLLMs (e.g., a 37.6% improvement on MM-SafetyBench (SD+OCR), and 71.3% on VLSafe for LLaVA-1.5-7B) without sacrificing their performance on utility.
@@ 2 Related Work
By integrating the capabilities of the visual perception with LLMs, MLLMs [22, 14, 3, 62, 1, 9] inherit the robust reasoning capabilities of LLMs and excel in dialogues incorporating with visual elements. Despite their impressive capabilities, it has been observed that SoTA MLLMs are increasingly vulnerable to malicious visual inputs [36]. Recent works can be bifurcated into two distinct approaches with respect to the injection of malicious content, with one line of works [20, 35] showing that embedding the malicious textual queries into images via typography can effectively circumvent the defense mechanisms of MLLMs, while the other [51, 15, 45, 57, 42, 2, 50, 4, 18] focusing on employing gradient-based techniques to create adversarial images that prompt generation of the harmful responses, revealing severe vulnerability.
@@
To enhance safety of MLLMs, a straightforward approach involves aligning MLLMs with specially-constructed red-teaming data [30, 68, 11]. However, red-teaming is labor-intensive and might not encompass all potential attack vectors. Another approach focuses on protecting MLLMs during inference. Wu et al. [61] introduce the manual crafting of system prompts delineating permissible and impermissible actions, which might become less effective as new attacks emerge. Wang et al. [58] employ safety steering vectors to adjust MLLM activation in response to unsafe inputs. However, this may overlook unsafe intents in images not detectable by text-centric safety vectors. The most relevant to ours are [44, 10]. Pi et al. [44] augments the MLLMs with an ancillary unsafe content detector and output detoxifier, which are external and necessitate the additional training on extensive datasets. Instead, our proposed ECSO solely leverages the intrinsic safety mechanism of the pre-aligned LLMs in MLLMs, and is devoid of any further training. Chen et al. [10] introduce a novel automatic self-moderation mechanism, enabling MLLMs to assess and adjust their responses against specific criteria. Despite the promising performance, we will demonstrate in Sec. 5.5.1 that MLLMs struggle to refine their responses when confronted with images, highlighting the limitation of [10].
@@ 3 Preliminary Observations
(a) Safety persists in MLLMs without images.
(b) MLLMs are aware of unsafe responses.
Figure 1: (a) Comparisons of harmless rate (%) of model responses with and without images on five state-of-the-art MLLMs. (b) Accuracy of MLLMs discrimination (with and without images) on whether their own responses are safe or not .
@@ 3.1 Safety Mechanism Persists in MLLM
In contrast to previous findings suggesting that MLLMs struggle to inherit the safety mechanisms in LLMs [68, 35, 44], here we present a more nuanced view that MLLMs can retain the safeguard when images are not shown to the MLLM. In the following, we experiment with the VLSafe dataset [11], which contains 1,110 pairs of queries and images. This dataset has two key features: (1) The malicious queries are paired with benign images; and (2) The images are auxiliary inputs in nature, i.e. the queries can be answered without referencing the images. These features allow us to dissect the interaction between visual features and safety mechanism by evaluating MLLMs’ responses with and without the input images.
Figure 0(a) compares the harmless rates of MLLMs’ responses with and without the presence of images. As can be seen, when images are present, the MLLMs (except Qwen-VL-Chat [3]) are vulnerable to malicious queries, demonstrating a mere 20% harmless rate. On the other hand, when images are removed from the queries, all models achieve nearly 100% harmless rate. We hypothesize that this discrepancy arises from a distribution shift caused by integrating LLMs with the visual inputs. Specifically, the incorporation of images alters the pre-aligned embedding space of LLMs, rendering existing defense mechanisms ineffective [44]. However, we will show in Sec. 4.2 that the safety mechanism can be reactivated once images undergo a query-aware I2T transformation.
@@ 3.2 MLLMs are Aware of Their Own Unsafe Responses
While MLLMs are susceptible to generating the harmful content, we investigate whether they are aware of their own safety issues. In this experiment, we collect 1000 responses from LLaVA-1.5-7B, in which 500 of them are safe and the remaining 500 are unsafe as evaluated by GPT-4  [^1] https://chatgpt.ust.hk (and double-checked manually) [^2] More detailed description on the dataset can be found in Appendix 0.A.3.. We then prompt the MLLMs (detailed in Sec. 4.1) to classify the responses as safe or unsafe. Figure 0(b) shows the classification accuracies obtained on five MLLMs. As can be seen, though the MLLMs may generate unsafe responses, they exhibit a high degree of safety awareness. Notably, both LLaVA-1.5-7B and ShareGPT4V-7B achieve over 95% accuracy in their assessment, and it is crucial to highlight that unlike Figure 0(a), the assessment here is robust whether input images are presented or not. Therefore, safety awareness of MLLMs is not compromised by the presence of images. If no otherwise stated, we conduct MLLM safety discrimination with images by default in the remaining parts.
In summary, a significant discrepancy is observed between MLLMs’ ability to generate safe content (Figure 0(a)) and their capacity for safety discrimination (Figure 0(b)). This divergence may be attributed to the inherent ease of discrimination tasks over generation tasks, a hypothesis supported by [7], or understood through analogies drawn from scalable oversight and complexity theory [48]. In Sec. 4, we will explore how leveraging these insights can effectively mitigate the safety concerns associated with MLLMs.
@@ 4 Methodology
In this section, we propose ECSO (Eyes Closed, Safety On), a novel training-free MLLM protection strategy by exploiting the two observations in Sec. 3. The following sections detail the main steps, with an overview given in Figure 2.
@@ 4.1 Harmful Content Detection
As in Figure 2, we first prompt the MLLM $F_{\boldsymbol{\theta}}$ (with parameter $\boldsymbol{\theta}$) to output response $\tilde{y}$ to a (benign or malicious) user query with image $v$ and query $x$:
@@
The response $\tilde{y}$ may not be safe as MLLMs are susceptible to attacks. Using a prompt template $P_{\text{det}}$ for harm detection (Figure 3, left) to wrap $x$ and $\tilde{y}$, we ask the MLLM to discriminate the safety of its own $\tilde{y}$ (Step 2 in Figure 2):
@@
where $s$ indicates the predicted safety of the model response. The near-GPT-4 performance on safety assessment as seen in Sec. 3.2 supports this introspection. If the model response $\tilde{y}$ passes this self-checking, it will be presented to the user. On the contrary, if $\tilde{y}$ detected unsafe, we propose to first transform the image into text (Step 3 in Figure 2) and then query MLLMs again without visual inputs (Step 4 in Figure 2), which will be detailed in Sec. 4.2 and 4.3, respectively.
@@ 4.2 Query-Aware Image-to-Text (I2T) Transformation
To restore the intrinsic safety mechanism of the pre-aligned LLMs in MLLMs, we propose to transform the input query image to plain text. Any malicious content in the image that might induce harmful responses are then either converted to text or completely left away from the remaining procedure.
However, there may be information loss in the image-to-text (I2T) conversion. To retain the image information to the greatest extent, we use a prompt template $P_{\text{trans}}$ (Figure 3, middle) that includes the original question. The MLLM is then prompted to generate the query-aware caption $c$:
@@ 4.3 Safe Response Generation Without Images
where $\mathrm{null}$ denotes an empty input (i.e., the query image $v$ is excluded), and $P_{\text{gen}}$ is the prompt for safe response generation without images (Figure 3, right). Since images have been removed, the MLLM degrades to a text-only LLM. This step is effective in safeguarding the MLLM because any malicious contents that induce harmful responses are exposed to the safe pre-aligned LLM. To further underscore the priority of safety, we include the words “HARMLESS and ETHICAL” in the prompt during inference, as shown in Figure 3 (right).
@@ 5 Experiments
In this section, we empirically evaluate the proposed ECSO method. First, we introduce experimental settings in Sec. 5.1 and 5.2. Then, we assess ECSO from the following perspectives: (i) How well can ECSO protect the existing MLLMs? (Sec. 5.3) (ii) Can ECSO maintain the utility of MLLMs? (Sec. 5.4) (iii) Can ECSO serve as a data engine to produce data for safety alignment? (Sec. 5.6). Besides, we ablate the effects of the key components of ECSO in Sec. 5.5.
@@ 5.1 Models and Evaluation Protocols
Five SoTA MLLMs are considered, including the LLaVA-1.5-7B [33], ShareGPT4V-7B  [9], mPLUG-OWL2-7B [62], Qwen-VL-Chat [3] and InternLM-XComposer [64]. To demonstrate the effectiveness of our ECSO, we compare with directly prompting MLLMs and responding to queries (denoted as Direct).
@@
The safety degree of a model’s responses is measured by the harmless rate [52, 7], which is defined as the proportion of safe responses over all the responses $D$ (i.e., $\text{HR}=\sum_{d\in D}\frac{I(d)}{\left|D\right|}$), where $I(\cdot)$ is the indicator function with $I(\cdot)=1$ if the response is harmless (as judged by prompting GPT-4 followed by manual verification), and $I(\cdot)=0$ otherwise. The detailed prompt for GPT-4 can be found in Figure 11.
@@ 5.3 Evaluation of Safety
Table 1 compares the harmless rates on MM-SafetyBench by directly prompting LLaVA-1.5-7B (Direct) and prompting via the proposed ECSO. As can be seen, ECSO greatly boosts the safety of LLaVA-1.5-7B. Specifically, on average, the proposed ECSO method improves LLaVA-1.5-7B’s harmless rate from 31.7% to 90.3% when queried with OCR images, and from 32.1% to 86.4% when queried with SD+OCR images. In particular, ECSO offers much bigger safety gains on OCR and SD+OCR compared to SD. This is because SD is less effective in attacking MLLMs (as can be seen from Table 1). As most SD responses obtained by a direct prompting of LLaVA-1.5-7B are already benign, the improvement by ECSO is smaller. It is interesting that the harmless rate of ECSO even surpasses Text-Only (i.e., the upper bound of ECSO). This can be explained by the inclusion of the keywords “HARMLESS and ETHICAL” in Sec. 4.3, instructing LLMs to pay more attention and respond in a safer way.
@@
Recall from Sec. 3.1 that the MLLMs can only achieve satisfactory harmless rates when images are excluded. Now, with ECSO, we can maintain the safety of MLLMs while retaining the information in the images. Hence, we conclude that the proposed ECSO method can effectively reactivate the safety mechanism within MLLMs even with the presence of images via query-aware I2T transformation and safe response generation without images.

### v3

@@ Abstract
Multimodal large language models (MLLMs) have shown impressive reasoning abilities. However, they are also more vulnerable to jailbreak attacks than their LLM predecessors. Although still capable of detecting the unsafe responses, we observe that safety mechanisms of the pre-aligned LLMs in MLLMs can be easily bypassed with the introduction of image features. To construct robust MLLMs, we propose ECSO (Eyes Closed, Safety On), a novel training-free protecting approach that exploits the inherent safety awareness of MLLMs, and generates safer responses via adaptively transforming unsafe images into texts to activate the intrinsic safety mechanism of pre-aligned LLMs in MLLMs. Experiments on five state-of-the-art (SoTA) MLLMs demonstrate that ECSO enhances model safety significantly (e.g., 37.6% improvement on the MM-SafetyBench (SD+OCR) and 71.3% on VLSafe with LLaVA-1.5-7B), while consistently maintaining utility results on common MLLM benchmarks. Furthermore, we show that ECSO can be used as a data engine to generate supervised-finetuning (SFT) data for MLLM alignment without extra human intervention.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) [14, 3, 66, 9, 23, 68] have attracted significant attention for their remarkable multimodal capabilities. Building upon the Large Language Models (LLMs) [58, 28, 56, 12], they are aligned with a pre-trained visual encoder using text-image datasets [34, 36, 33], empowering LLMs to conduct conversations with image inputs. Despite these accomplishments, MLLMs encounter challenges in inheriting the safety mechanism of their LLM predecessors. In particular, though MLLMs are built upon LLMs that have been well-aligned with human morals and values [12, 59], they can be easily induced to generate unethical content with the introduction of image inputs [73, 38, 48].
To protect MLLMs, one can repeat training-based alignment strategies of LLMs on MLLMs, such as Supervised Finetuning (SFT) [63, 35, 42, 7] and Reinforcement Learning from Human Feedback (RLHF) [47, 13, 50]. However, this requires meticulous design of red-teaming queries to induce LLMs to generate harmful responses, and can become even more challenging when image inputs involved [48, 73]. Thus, the question is: “How can we transfer the pre-aligned safety mechanisms of LLMs to MLLMs?”
In this paper, we start by conducting a throughout analysis on the safety assessment ability of MLLMs. We observe that despite their susceptibility to malicious queries, (i) MLLMs exhibit clear awareness of unsafe content in their own responses [7]. (ii) The safety mechanism of pre-aligned LLMs persists in MLLMs, but is “suppressed” by the introduction of image features. However, this can be restored by simply removing the images. Building upon these insights, we propose ECSO (Eyes Closed, Safety On), a novel training-free MLLM protection strategy exploiting the intrinsic safety mechanisms of pre-aligned LLMs. When presented with an image input with a user query, ECSO first leverages the safety awareness of MLLMs to assess safety of their own responses in a post-hoc manner. Once unsafe initial responses are detected, ECSO converts the image inputs into texts via a query-aware image-to-text (I2T) transformation, and reduces MLLMs to (text-only) LLMs. Safe response generation without images is then performed to restore the safety mechanism of pre-aligned LLMs. Experiments on five MLLM benchmarks demonstrate that the proposed ECSO can significantly enhance model safety without sacrificing the utility performance on common MLLM benchmarks. Moreover, we show that ECSO can be used as a data engine for the generation of SFT data to align MLLMs without extra human intervention.
@@
    We demonstrate that MLLMs, though susceptible to jailbreaking attacks, can detect unsafe content in their own responses and also inherit the safety mechanisms from pre-aligned LLMs that have been inadvertently suppressed.
@@
    ECSO significantly enhances the safety of five SoTA MLLMs, without sacrificing their performance on utility.
@@ 2 Related Work
By integrating the capabilities of visual perception with LLMs, MLLMs [23, 14, 3, 66, 1, 9] inherit robust reasoning capabilities of LLMs and excel in dialogues incorporating with visual elements. Despite their impressive capabilities, it has been observed that SoTA MLLMs are increasingly vulnerable to malicious visual inputs [39]. Recent works can be bifurcated into two approaches with respect to the injection of malicious content. One line of works [21, 38] show that embedding the malicious textual queries into images via typography can effectively circumvent the defense mechanisms of MLLMs. The other approach [70, 54, 15, 49, 60, 46, 2, 53, 4, 18] focus on employing gradient-based techniques to create adversarial images that prompt generation of the harmful responses, revealing severe vulnerability.
@@
To enhance safety of MLLMs, a straightforward approach involves aligning MLLMs with specially-constructed red-teaming data [31, 73, 11]. However, red-teaming is labor-intensive and may not encompass all potential attack vectors. Another approach focuses on protecting MLLMs during inference. Wu et al. [65] introduce the manual crafting of system prompts delineating permissible and impermissible actions. However, this may become less effective when new attacks emerge. Wang et al. [61] employ safety steering vectors to adjust MLLM activation in response to the unsafe inputs. However, this may overlook unsafe intents in images that are not detectable by text-centric safety vectors. Most relevant to ours are the works in [10, 48]. Chen et al. [10] introduce a novel automatic self-moderation mechanism, enabling MLLMs to assess and adjust their responses against specific criteria. Despite its promising performance, we will show in Sec. 5.5.1 that, even though instructed to respond safely, MLLMs still struggle to give responses when confronted with images, highlighting the limitation of [10]. Pi et al. [48] augments MLLMs with an ancillary unsafe content detector and output detoxifier, which are external and necessitate additional training on extensive datasets. Instead, the proposed ECSO solely leverages the intrinsic safety mechanism of the pre-aligned LLMs in MLLMs, and is devoid of any further training.
@@ 3 Preliminary Observations
Figure 1: (left) MLLMs are vulnerable to malicious questions when queried with images but can restore safety when images are excluded. (right) Comparisons of harmless rate (%) of model responses with and without images on five state-of-the-art MLLMs.
Figure 2: (left) Though vulnerable to malicious questions, MLLMs are aware of the unsafe responses of their own. (right) Accuracy of MLLMs discrimination (with and without images) on whether their own responses are safe or not .
@@ 3.1 Safety Mechanism Persists in MLLM
In contrast to previous findings suggesting that MLLMs struggle to inherit the safety mechanisms in LLMs [73, 38, 48], here we present a more nuanced view that MLLMs can retain the safeguard when images are not shown to the MLLM. In the following, we perform experiment on the VLSafe dataset [11], which contains 1,110 pairs of queries and images. This dataset has two key features: (1) The malicious queries are paired with benign images; and (2) The input images are auxiliary, i.e. the queries can be answered without referencing the images. These features allow us to dissect the interaction between visual features and safety mechanism by evaluating MLLMs’ responses with and without input images.
Figure 1 compares the harmless rates of MLLMs’ responses with and without the presence of images. As can be seen, when images are present, the MLLMs (except Qwen-VL-Chat [3]) are vulnerable to malicious queries, demonstrating a mere 20% harmless rate. On the other hand, when images are removed from the queries, all models achieve nearly 100% harmless rate. We hypothesize that this discrepancy arises from a distribution shift caused by integrating LLMs with the visual inputs. Specifically, the incorporation of images alters the pre-aligned embedding space of LLMs, rendering existing defense mechanisms ineffective [48]. However, we will show in Sec. 4.2 that the safety mechanism can be reactivated once images undergo a query-aware I2T transformation.
@@ 3.2 MLLMs are Aware of Their Own Unsafe Responses
While MLLMs are susceptible to generating harmful content, we investigate whether they are aware of their own safety issues. In this experiment, we collect 1000 responses from LLaVA-1.5-7B. 500 of them are safe and the remaining 500 are unsafe as evaluated by GPT-4111https://chatgpt.ust.hk and double-checked manually. [^2] More detailed description on the dataset can be found in Appendix 0.A.3. We then prompt the MLLMs (detailed in Sec. 4.1) to classify the responses as safe or unsafe. Figure 2 shows the classification accuracies obtained on five MLLMs. As can be seen, though the MLLMs may generate unsafe responses, they exhibit a high degree of safety awareness. Notably, both LLaVA-1.5-7B and ShareGPT4V-7B achieve over 95% accuracy in their assessment. It is crucial to highlight that unlike Figure 1, the assessment here is robust whether the input images are presented or not. Therefore, safety awareness of MLLMs is not compromised by the presence of images. If not otherwise stated, we conduct MLLM safety discrimination with images by default in the sequel.
In summary, a significant discrepancy is observed between MLLMs’ ability to generate safe content (Figure 1) and their capacity for safety discrimination (Figure 2). This divergence may be attributed to the inherent ease of discrimination tasks over generation tasks (a hypothesis supported by [7]) or understood through analogies drawn from scalable oversight and complexity theory [52]. In Sec. 4, we will explore how leveraging these insights can effectively mitigate the safety concerns associated with MLLMs.
@@ 4 Methodology
In this section, we propose ECSO (Eyes Closed, Safety On), a novel training-free MLLM protection strategy by exploiting the two observations in Sec. 3. An overview of the main steps is shown in Figure 3.
@@ 4.1 Harmful Content Detection
Given a (benign or malicious) user query with image $v$ and query $x$, we first prompt the MLLM $F_{\boldsymbol{\theta}}$ (with parameter $\boldsymbol{\theta}$) to output response $\tilde{y}$ (Figure 3):
@@
The response $\tilde{y}$ may not be safe as MLLMs are susceptible to attacks. Using a prompt template $P_{\text{det}}$ for harm detection (Figure 4, left) to wrap $x$ and $\tilde{y}$, we ask the MLLM to discriminate the safety of its own $\tilde{y}$ (Step 2 in Figure 3), and output the predicted safety $s$ of the model response:
@@
The near-GPT-4 performance on safety assessment, as seen in Sec. 3.2, supports this introspection. If the model response $\tilde{y}$ passes this self-checking, it will be presented to the user. On the contrary, if $\tilde{y}$ is detected as unsafe, we propose to first transform the image into text (Step 3 in Figure 3) and then query the MLLMs again without visual inputs (Step 4 in Figure 3). These will be detailed in Sec. 4.2 and 4.3, respectively.
@@ 4.2 Query-Aware Image-to-Text (I2T) Transformation
To restore the intrinsic safety mechanism of the pre-aligned LLMs in MLLMs, we propose to transform the input query image to plain text. Any malicious content in the image that might induce harmful responses are then either converted to text or completely left away from the remaining procedure. However, there may be information loss in the image-to-text (I2T) conversion. To retain the image information to the greatest extent, we use a prompt template $P_{\text{trans}}$ (Figure 4, middle) that includes the original question. The MLLM is then prompted to generate the query-aware caption $c$:
@@ 4.3 Safe Response Generation Without Images
where $\mathrm{null}$ denotes an empty input (i.e., the query image $v$ is excluded), and $P_{\text{gen}}$ is the prompt for safe response generation without images (Figure 4, right). Since images have been removed, the MLLM reduces to a text-only LLM. This step is effective in safeguarding the MLLM because any malicious contents that induce harmful responses are exposed to the safe pre-aligned LLM. To further underscore the priority of safety, we include the words “HARMLESS and ETHICAL” in the prompt during inference, as shown in Figure 4 (right).
@@ 5 Experiments
In this section, we empirically evaluate the proposed ECSO. First, we introduce the experimental settings in Sec. 5.1 and 5.2. Then, we assess ECSO from the following perspectives: (i) How well can ECSO protect the existing MLLMs? (Sec. 5.3) (ii) Can ECSO maintain the utility of MLLMs? (Sec. 5.4) (iii) Can ECSO serve as a data engine to produce data for safety alignment? (Sec. 5.6). Besides, we ablate the effects of the key components of ECSO in Sec. 5.5.
@@ 5.1 Models and Evaluation Protocols
Five SOTA MLLMs are considered, including the LLaVA-1.5-7B [36], ShareGPT4V-7B  [9], mPLUG-OWL2-7B [66], Qwen-VL-Chat [3] and InternLM-XComposer [68]. To demonstrate the effectiveness of the proposed ECSO, we compare with directly prompting the MLLMs and responding to queries (denoted Direct).
@@
The safety degree of a model’s responses is measured by the harmless rate [55, 7], which is defined as the proportion of safe responses over all the responses $D$ (i.e., $\text{HR}=\sum_{d\in D}\frac{I(d)}{\left|D\right|}$, where $I(\cdot)$ is the indicator function with $I(\cdot)=1$ if the response is harmless (as judged by prompting GPT-4 followed by manual verification), and $I(\cdot)=0$ otherwise). The detailed prompt for GPT-4 can be found in Figure 11.
@@ 5.3 Evaluation of Safety
Table 1 compares the harmless rates on MM-SafetyBench by directly prompting LLaVA-1.5-7B (Direct) and prompting via the proposed ECSO. As can be seen, ECSO greatly boosts the safety of LLaVA-1.5-7B. Specifically, on average, the proposed ECSO improves LLaVA-1.5-7B’s harmless rate from 31.7% to 90.3% when queried with OCR images, and from 32.1% to 86.4% when queried with SD+OCR images. In particular, ECSO offers much bigger safety gains on OCR and SD+OCR compared to SD. This is because SD is less effective in attacking MLLMs (as can be seen from Table 1). As most SD responses obtained by a direct prompting of LLaVA-1.5-7B are already benign, the improvement by ECSO is smaller. It is interesting that the harmless rate of ECSO even surpasses Text-Only (i.e., the upper bound of ECSO). This can be explained by the inclusion of the keywords “HARMLESS and ETHICAL” in Sec. 4.3, instructing LLMs to pay more attention and respond in a safer way.
@@
Recall from Sec. 3.1 that the MLLMs can only achieve satisfactory harmless rates when images are excluded. Now, with ECSO, we can maintain the safety of MLLMs while retaining the information in the images. Hence, we conclude that the proposed ECSO can effectively reactivate the safety mechanism within MLLMs even with the presence of images via query-aware I2T transformation and safe response generation without images.